# Dispersive regime of multiphoton qubit-oscillator interactions

### v1

@@ VI Summary and Conclusions
Typical SW-type perturbation theory transformations are used to decouple low-energy and high-energy subspaces; this allows us to obtain an effective description of a complicated many-body system within its low-energy subspaces [53]. The presented generalization provides an accurate description for the lower-energy states, and the accuracy becomes ever-vanishing for higher-energy states, as expected from a low-energy effective description of the system. In the usual linear interaction qubit-oscillator dispersive regime, the accuracy is commonly bounded by a critical photon number, $n_{\text{c}}=g^{2}/4\Delta^{2}$, which signifies the growing deviation between the perturbative and full system Hamiltonians for oscillator photon occupation numbers beyond it. A general bound on the spectra to match this new notion of critical photon number is difficult to obtain for the multiphoton dispersive regime as it depends on multiple parameters such as $n$, $g_{n}$, and $\Delta_{n}$. In practice, such a bounding parameter will depend on the additional imperfections and spurious effects of a given implementation choice.
A surprising find is the non-RWA approximation is less accurate than its RWA counterpoart. This suggests that the role of the counter-rotating terms is not accounted for by a simple extension of the SW transformation generators. Unlike the dispersive regime of the linear interaction, we find the multiphoton non-RWA dispersive regime to not provide accurate descriptions of the system at larger coupling strengths.
In the generalized case of multiple qubits nonlinearly coupled to a shared oscillator, we found similar trends to the single qubit-oscillator case with some differences due to the cooperative effects. Generally, the perturbative spectra presented provide an accurate description for the multiqubit case when the qubits’ couplings are in the individual qubit-oscillator regime of validity. When one of the qubits is outside its perturbative regime of validity, the perturbative descriptions are applicable for a much smaller range of parameters. The second key difference is the emergence of a photon-number-dependent qubit-qubit interaction. Incidentally, this allows for an effective tuning of the two-qubit system parameters . The state of the oscillator can then be used to tune the effective multiqubit system parameters.
@@ Acknowledgements.
M.A. was supported by the Institute for Quantum Computing (IQC) through funding provided by Transformative Quantum Technologies (TQT). S.A. was supported by Japan’s Ministry of Education, Culture, Sports, Science and Technology’s Quantum Leap Flagship Program Grant No. JPMXS0120319794.
@@ Appendix A Combinatorial aspects of multiphoton perturbation theory
We may now find the second order RWA and non-RWA dispersive Hamiltonians presented in the main text. The BCH expansion truncated at second order for the RWA Hamiltonian reads
$\hat{H}_{\text{Disp,RWA}}^{(n),2}$ $=\hat{H}_{n-\text{JC}}+\lambda_{n}[\hat{H}_{n-\text{JC}},\hat{X}_{-n}]$
$\,\,\,\,\,\,\,+\frac{\lambda_{n}^{2}}{2}[[\hat{H}_{n-\text{JC}},\hat{X}_{-n}],\hat{X}_{-n}]$
$=\hat{H}_{n-\text{JC}}-\Delta_{n}\lambda_{n}\hat{X}_{-n}$
@@
where we used $[\hat{H}_{\text{n-JC}},\hat{X}_{-n}]=-\Delta_{n}\hat{X}_{-n}+g_{n}[\hat{X}_{+n},\hat{X}_{-n}].$ The last term is on the order of $\lambda_{n}^{3}$ and, thus, neglected from Hamiltonian of Eq. (III.1) presented in the main text. A similar calculation for the non-RWA Hamiltonian yields
$\hat{H}_{\text{Disp}}^{(n),2}$ $=\hat{H}_{n-\text{R}}+\lambda_{n}[\hat{H}_{\text{n-R}},\hat{X}_{-n}]+\overline{\lambda}_{n}[\hat{H}_{\text{n-R}},\hat{Y}_{-n}]+\frac{\lambda_{n}^{2}}{2}[[\hat{H}_{n-\text{R}},\hat{X}_{-n}],\hat{X}_{-n}]+\frac{\overline{\lambda}_{n}^{2}}{2}[[\hat{H}_{n-\text{R}},\hat{Y}_{-n}],\hat{Y}_{-n}]$
$\,\,\,\,\,\,\,+\frac{\lambda_{n}\overline{\lambda}_{n}}{2}[[\hat{H}_{n-\text{R}},\hat{X}_{-n}],\hat{Y}_{-n}]+\frac{\overline{\lambda}_{n}\lambda_{n}}{2}[[\hat{H}_{n-\text{R}},\hat{Y}_{-n}],\hat{X}_{-n}]$
$=\hat{H}_{n-\text{R}}-\Delta_{n}\lambda_{n}\hat{X}_{-n}+\left(g_{n}\lambda_{n}-\frac{\Sigma_{n}\lambda_{n}\overline{\lambda}_{n}}{2}\right)[\hat{Y}_{+n},\hat{X}_{-n}]+\left(g_{n}\lambda_{n}-\frac{-\Delta_{n}\lambda_{n}^{2}}{2}\right)[\hat{X}_{+n},\hat{X}_{-n}]-\Sigma_{n}\overline{\lambda}_{n}\hat{Y}_{-n}$
@@
where we used $[\hat{H}_{\text{n-R}},\hat{X}_{-n}]=-\Delta_{n}\hat{X}_{+n}+g_{n}[\hat{Y}_{+n},\hat{X}_{-n}]+g_{n}[\hat{X}_{+n},\hat{X}_{-n}]$ and $[\hat{H}_{\text{n-R}},\hat{X}_{-n}]=-\Sigma_{n}\hat{Y}_{+n}+g_{n}[\hat{Y}_{+n},\hat{Y}_{-n}]+g_{n}[\hat{X}_{+n},\hat{Y}_{-n}].$ The last eight terms comprising nested commutators are of third order in $\lambda_{n}$ and $\overline{\lambda}_{n}$, and, thus, are not included in the non-RWA Hamiltonian of Eq. (III.2).

### v2

@@ VI Summary and Conclusions
Typical SW-type perturbation theory transformations are used to decouple low-energy and high-energy subspaces; this allows us to obtain an effective description of a complicated many-body system within its low-energy subspaces [57]. The presented generalization provides an accurate description for the lower-energy states, and the accuracy becomes ever-vanishing for higher-energy states, as expected from a low-energy effective description of the system. In the usual linear interaction qubit-oscillator RWA dispersive regime, the accuracy is commonly bounded by a critical photon number, $n_{\text{ph,c}}=\Delta^{2}/4g^{2}$, which signifies the growing deviation between the perturbative and full system Hamiltonians for higher photon occupation numbers. We derived a generalized heuristic critical photon number for the multiphoton RWA dispersive regime, $n_{\text{ph},c}^{(n)}\propto(\Delta_{n}/g_{n})^{2/n},$ that defines a narrowing regime of validity for higher-energy states. Additionally, as seen in Fig. 2, the regime of validity shrinks for higher order $n.$ In the RWA regime, we found that our analytic spectra were in excellent agreement with the numerical results. In the non-RWA regime ($|\Delta_{n}|\sim\Sigma_{n}$), the RWA analytic spectra become less accurate of an approximation, and we resorted to an extension of the SW transformation that takes the counter-rotating interaction terms, $\hat{\sigma}_{-}\hat{a}^{\dagger n}+\hat{\sigma}_{+}\hat{a}^{n}$, into account. In this regime, the non-RWA spectra are an excellent match for the numerical spectra.
In the generalized case of multiple qubits nonlinearly coupled to a shared oscillator, we found similar trends to the single qubit-oscillator case with some differences due to the cooperative effects. Generally, the perturbative spectra presented provide an accurate description for the multiqubit case when the qubits’ collective cooperative effects are weak. A key difference is that in the presence of stronger cooperative effects, i.e. stronger individual qubit coupling strengths, the analytic RWA and non-RWA spectra deviate at smaller coupling strengths in their respective regimes of validty. The second key difference is the emergence of a photon-number-dependent qubit-qubit interaction. Incidentally, this allows for an effective tuning of the two-qubit system parameters . The state of the oscillator can then be used to tune the effective multiqubit system parameters.
@@ Acknowledgements.
We thank the anonymous referees for their useful comments that significantly helped improve the manuscript. M.A. was supported by the Institute for Quantum Computing (IQC) through funding provided by Transformative Quantum Technologies (TQT). S.A. was supported by Japan’s Ministry of Education, Culture, Sports, Science and Technology’s Quantum Leap Flagship Program Grant No. JPMXS0120319794.
@@ Appendix A Combinatorial aspects of multiphoton perturbation theory
We may now find the second-order RWA and non-RWA dispersive Hamiltonians presented in the main text. The BCH expansion truncated at second order for the RWA Hamiltonian reads
$\hat{H}_{\text{Disp,RWA}}^{(n),2}$ $=\hat{H}_{n\text{JC}}+\lambda_{n}[\hat{H}_{n\text{JC}},\hat{X}_{-n}]$
$\,\,\,\,\,\,\,+\frac{\lambda_{n}^{2}}{2}[[\hat{H}_{n\text{JC}},\hat{X}_{-n}],\hat{X}_{-n}]$
$=\hat{H}_{n\text{JC}}-\Delta_{n}\lambda_{n}\hat{X}_{-n}$
@@
where we used $[\hat{H}_{n\text{JC}},\hat{X}_{-n}]=-\Delta_{n}\hat{X}_{-n}+g_{n}[\hat{X}_{+n},\hat{X}_{-n}].$ The last term is on the order of $\lambda_{n}^{3}$ and, thus, neglected from the Hamiltonian of Eq. (III.1) presented in the main text. A similar calculation for the non-RWA Hamiltonian yields
$\hat{H}_{\text{Disp}}^{(n),2}$ $=\hat{H}_{n\text{R}}+\lambda_{n}[\hat{H}_{\text{n-R}},\hat{X}_{-n}]+\overline{\lambda}_{n}[\hat{H}_{\text{n-R}},\hat{Y}_{-n}]+\frac{\lambda_{n}^{2}}{2}[[\hat{H}_{n\text{R}},\hat{X}_{-n}],\hat{X}_{-n}]+\frac{\overline{\lambda}_{n}^{2}}{2}[[\hat{H}_{n\text{R}},\hat{Y}_{-n}],\hat{Y}_{-n}]$
$\,\,\,\,\,\,\,+\frac{\lambda_{n}\overline{\lambda}_{n}}{2}[[\hat{H}_{n\text{R}},\hat{X}_{-n}],\hat{Y}_{-n}]+\frac{\overline{\lambda}_{n}\lambda_{n}}{2}[[\hat{H}_{n\text{R}},\hat{Y}_{-n}],\hat{X}_{-n}]$
$=\hat{H}_{n\text{R}}-\Delta_{n}\lambda_{n}\hat{X}_{-n}+\left(g_{n}\lambda_{n}-\frac{\Sigma_{n}\lambda_{n}\overline{\lambda}_{n}}{2}\right)[\hat{Y}_{+n},\hat{X}_{-n}]+\left(g_{n}\lambda_{n}-\frac{-\Delta_{n}\lambda_{n}^{2}}{2}\right)[\hat{X}_{+n},\hat{X}_{-n}]-\Sigma_{n}\overline{\lambda}_{n}\hat{Y}_{-n}$
@@
where we used $[\hat{H}_{\text{n-R}},\hat{X}_{-n}]=-\Delta_{n}\hat{X}_{+n}+g_{n}[\hat{Y}_{+n},\hat{X}_{-n}]+g_{n}[\hat{X}_{+n},\hat{X}_{-n}]$ and $[\hat{H}_{\text{n-R}},\hat{Y}_{-n}]=-\Sigma_{n}\hat{Y}_{+n}+g_{n}[\hat{Y}_{+n},\hat{Y}_{-n}]+g_{n}[\hat{X}_{+n},\hat{Y}_{-n}].$ The last eight terms comprising nested commutators are of third order in $\lambda_{n}$ and $\overline{\lambda}_{n}$, and, thus, are not included in the non-RWA Hamiltonian of Eq. (III.2).
Appendix B Stabilization of multiphoton qubit-oscillator interaction models
Figure 7: Schematic diagram that illustrates the stabilization of the Hamiltonian. The gray dashed line represents the unperturbed harmonic oscillator potential. The red line shows the unstabilized potential with harmonic, $x^{2}$, and cubic, $x^{3}$, contributions. In the unstabilized case, we obtain a continuous spectrum corresponding to unbounded states, with energies extending to $\infty$- ∞. The blue line is the stabilized potential where a small positive quartic term, $x^{4}$, is added to the unstabilized potential to ensure that it is bounded from below. The infinite negative energy states in the unstabilized case are replaced by bounded well-behaved states in the stabilized potential. The horizontal red dotted line represents the energy barrier that helps isolate the metastable states in the shallow well and make them largely insensitive to the details of the potential in the deep well.
Figure 8: Spectrum of the three-photon Rabi model. (a) Unstabilized spectrum of the three-photon Rabi Hamiltonian [Eq. (III) with $\eta=0$] with an oscillator Hilbert space truncation of $N_{T}=300$. (b) Stabilized spectrum with $\eta=0.02$ and truncation size $N_{T}=8000.$ (c) Filtered spectrum in which we take the spectrum in (b) and keep only states with an average photon number less than 20.
In the main text, we discussed the serious stability issues with the $n$-photon Rabi model. For $n\geq 3$, the case $g=0$ is an isolated point where the spectrum is discrete and energetically stable. The spectrum becomes continuous and unbounded from below for any $g\neq 0$. In this appendix, we discuss how higher-order stabilizing terms arise in physical implementations of the n-photon Rabi model, using a realistic superconducting circuit implementation as an example. Then, we present some numerical results that demonstrate how the stabilization works and why our perturbative approach is justified. It is worth noting here that a recent study used a similar idea, in which the two-photon quantum Rabi model in the unstable parameter regime was stabilized by adding an $\left(\hat{a}+\hat{a}^{\dagger}\right)^{4}$ interaction term in the Hamiltonian [70].
B.1 Stabilizing terms in physical implementations
We consider a superconducting circuit implementation of an effective $n$-photon qubit-oscillator Hamiltonian. One specific example is the circuit proposed in Ref. [22], where a transmon qubit is coupled to an LC oscillator by means of a direct-current superconducting quantum interference device (dc SQUD). The transmon and oscillator are represented by their conjugate phase and charge variables, $\hat{\varphi}_{k}$ and $\hat{n}_{k}$, that satisfy the usual commutation relation $[\hat{\varphi}_{k},\hat{n}_{k}]=i$ with $k=t$ for the transmon and $k=o$ for the oscillator, and $[\hat{\varphi}_{t},\hat{n}_{o}]=[\hat{\varphi}_{o},\hat{n}_{t}]=0$. These conjugate variables can be written in terms of creation and annihilation operators as
$\hat{\varphi}_{t}=\varphi_{t}(\hat{b}^{\dagger}+\hat{b}),$ (64)
$\hat{n}_{t}=in_{t}(\hat{b}^{\dagger}-\hat{b}),$ (65)
$\hat{\varphi}_{o}=\varphi_{o}(\hat{a}^{\dagger}+\hat{a}),$ (66)
and
$\hat{n}_{o}=in_{o}(\hat{a}^{\dagger}-\hat{a}),$ (67)
where $\varphi_{k}$ and $n_{k}$ are the zero-point fluctuation amplitudes in the phase and charge for the respective systems. In theory, an asymmetric dc SQUID mediates all possible combinations of nonlinear interaction terms between the transmon and oscillator, i.e. terms that take the form $\hat{\varphi}_{t}^{m}\hat{\varphi}_{o}^{n}$ for nonnegative integers $m$ and $n$, all occurring together with varying effects depending on the coupling strengths and resonance conditions. For our purposes, we typically seek to implement and utilize only one interaction term of the form $\hat{\varphi}_{t}\hat{\varphi}_{o}^{n}$. Under the two-level approximation, we have $\hat{\varphi}_{t}\propto(\hat{b}^{\dagger}+\hat{b})\simeq\hat{\sigma}_{+}+\hat{\sigma}_{-}$, which effectively leads to an interaction Hamiltonian $\propto(\hat{\sigma}_{+}+\hat{\sigma}_{-})(\hat{a}^{\dagger}+\hat{a})^{n}$. The interaction term $(\hat{\sigma}_{+}\hat{a}^{\dagger n}+\hat{\sigma}_{-}\hat{a}^{n})$ is an approximation obtained under additional weak coupling and specific resonance conditions. To arrive at such an interaction Hamiltonian, we assume that all other interactions (e.g. those of the form $\hat{\varphi}_{t}^{m}\hat{\varphi}_{o}^{l}$ with $l\neq n$) have negligible effects on the spectrum and dynamics. A standard approximation in this context is that of low photon numbers and short evolution times. When one considers states with higher photon numbers and/or dynamics at long timescales, the effects of other interactions are generally no longer negligible. Hence, one must always keep in mind that a realistic system, as described here, has higher-order terms that invalidate the model when the photon number becomes large. The exact way in which the higher-order terms alter the behavior of the system depends on the specifics of the physical system. In this appendix, we consider a stabilizing term of the form $\hat{\varphi}_{o}^{m}$ with $m>n$.
B.2 Intuitive picture of spectrum stabilization
To give an intuitive picture of how the stabilization of the spectrum works, it is convenient to consider the $n$-photon full Rabi Hamiltonian stabilized by an $m$th-order term
$\hat{H}_{n\text{fR}}=$ $\omega_{o}\hat{a}^{\dagger}\hat{a}+\frac{\omega_{q}}{2}\hat{\sigma}_{z}+g_{n}\hat{\sigma}_{x}(\hat{a}+\hat{a}^{\dagger})^{n}$
$+\eta g_{n}(\hat{a}+\hat{a}^{\dagger})^{m},$ (68)
with $m>n$ and $g_{n},\eta>0$. In the regime $\omega_{q}\ll\omega_{o}$, a good approximation for the spectrum is obtained by focusing on one of the two cases $\hat{\sigma}_{x}=\pm 1$ [71]. If we take $n=3$ and $m=4$, and we set $\hat{\sigma}_{x}=-1$, we obtain the Hamiltonian
$\hat{H}_{n\text{fR}}\Big{|}_{\sigma_{x}=-1}=$ $-\frac{\omega_{o}}{4}(\hat{a}^{\dagger}-\hat{a})^{2}+\frac{\omega_{o}}{4}(\hat{a}+\hat{a}^{\dagger})^{2}$
$-g_{3}(\hat{a}+\hat{a}^{\dagger})^{3}+\eta g_{3}(\hat{a}+\hat{a}^{\dagger})^{4},$ (69)
up to a constant. The Hamiltonian in Eq. (B.2) describes a single-particle problem in which $i(\hat{a}^{\dagger}-\hat{a})$ plays the role of the momentum variable $p$ while $(\hat{a}^{\dagger}+\hat{a})$ plays the role of the position variable $x$. The first term represents the kinetic energy, while the last three terms represent the potential energy. The potential energy of this Hamiltonian is illustrated in Fig. 7. If we ignore the last term, we obtain a function that is unbounded from below. The potential energy can be made infinite and negative by going to large values of $x$. Furthermore, the spectrum of such an unbounded Hamiltonian is continuous; there are propagating wave states that correspond to all energies from $\infty$- ∞ to +$\infty$+ ∞. The stabilizing term ensures that this situation does not occur. For sufficiently large $x$, the $x^{4}$ term will always be the dominant term in the potential energy. As a result, the Hamiltonian is bounded from below and the spectrum is discrete. There will be negative energy solutions. The smaller the value of the stabilizing term coefficient $\eta$, the larger (in absolute value) these energies. However, the most serious consequences of having a Hamiltonian that is unbounded from below are eliminated.
Perhaps more importantly, the potential energy has a local minimum around $x=0$. There is an energy barrier that separates the local minimum (shallow well) from the global minimum (deep well). As a result, one can expect that the energy levels in the shallow well that are well below the barrier height will still exist and only slightly be affected by the existence of a deep well relatively far away in $x$. Furthermore, for small values of $\eta$, the stabilizing term is negligible at small values of $x$. As a result, one can expect a perturbative treatment that ignores the stabilizing term to produce accurate results at small values of $x$, even if the spectrum as a whole exhibits singular behavior in this case.
A final point that is worth mentioning here is that the quantum states in the shallow well can be considered metastable states. As long as the barrier is high, the energy levels are close to those of the unperturbed Hamiltonian. As the coupling strength $g$ increases and the energy barrier decreases, the states in the shallow well become increasingly prone to tunneling into the deep well, and the corresponding energy levels gradually disappear from the spectrum. A simple analysis of the potential energy function in Eq. (B.2) shows that the disappearance of the lowest energy levels in the shallow well occurs when $g_{3}\sim\eta\omega_{o}$, which is an expression specific to our definition of the coefficient in the stabilizing term. This behavior will be observed in the spectra shown in the next subsection of this appendix (up to a minor quantitative difference, considering that the next subsection uses a different Hamiltonian). Even in the absence of the stabilizing term ($\eta=0$), one could say that these metastable states of the Hamiltonian exist and that they have a finite lifetime to tunnel into the continuum, after which the system relaxes into a qualitatively different state that is not described by the model Hamiltonian $\hat{H}_{n\text{fR}}$.
B.3 Numerical results on stabilization
For purposes of analyzing the spectrum of $\hat{H}_{n\rm JC}$ in the low-photon-number regime, it is convenient to consider a stabilizing term of the form $a^{\dagger m/2}a^{m/2}$ which is obtained from an oscillator phase term $\hat{\varphi}_{o}^{m}$. It turns out that the stabilizing term $a^{\dagger m/2}a^{m/2}$ is more effective in stabilizing the Hamiltonian than the term $(\hat{a}+\hat{a}^{\dagger})^{m}$. The reason is that there are states with large photon numbers but small $(\hat{a}+\hat{a}^{\dagger})$, which makes it possible for some states to have large values of $(\hat{a}^{n}+\hat{a}^{\dagger n})$ but small values of $(\hat{a}+\hat{a}^{\dagger})^{m}$.
In Fig. 8, we compare energy levels obtained using the unstabilized Hamiltonian with those obtained using the stabilized Hamiltonian. The vast majority of the energy levels behave differently in the two cases. Furthermore, the energy levels around zero energy (and negative energies) never converge if we increase the number of photons (i.e. the Hilbert space size) in the simulation. In contrast, the small-energy part of the spectrum of the stabilized Hamiltonian is independent of simulation size, provided a sufficient Hilbert space size is used. Setting aside this part of the spectrum and focusing on the small-slope energy levels in the weak-coupling regime, we find very good agreement between the spectra of the stabilized and unstabilized Hamiltonians. These energy levels correspond to the metastable states at low photon numbers described in the previous subsection.
As a further illustration of the behavior of these metastable states, in Fig. 8(c) we plot a spectrum in which we keep only the energy levels that correspond to eigenstates with $\left\langle\hat{a}^{\dagger}\hat{a}\right\rangle<20$, i.e. states with fewer than 20 photons on average. This way we keep only states that are localized near $x=0$. As expected, all the steep lines in the spectrum are eliminated, because they correspond to states that traverse a large range of $x$ values. Only the states that are confined in the shallow well remain in this spectrum. these energy levels are the ones for which we can expect the perturbative treatment to be valid. Furthermore, as the coupling strength increases and starts to approach $\eta\omega_{o}$, most metastable states disappear in rapid succession, as expected.
We finally note that we did similar simulations for $\omega_{q}=3.5\omega_{o}$ and found that, while the details of the weak-coupling spectrum change in accordance with the dispersive regime formulas in the main text, the point at which the states become unstable and disappear from the spectrum is almost the same as in the case $\omega_{q}=3.1\omega_{o}$, i.e. as shown in Fig. 8(c). Indeed, the dispersive regime energy shifts and the stability of states in the shallow well of the effective trapping potential are two separate issues.